# OptiMUS-0.3: Using large language models to model and solve optimization problems at scale

### v2

@@
Large Language Models, Optimization Modeling, AI Agents, Mixed Linear Integer Programming, Optimization Tools, Human-Computer Interaction
@@ 1 Introduction
  • We have developed a publicly available web application that allows practitioners to start using OptiMUS. This webapp also illustrates an initial approach to human computer interaction for LLM-powered optimization modeling. Notably, the interface makes it easy for the human to audit and correct the LLM at each stage of modeling, and allows the LLM to communicate uncertainty to the user in order to solicit corrections when needed.
An initial version of this work was published in a conference proceeding (Anonymous 2024) that introduced a smaller initial version of the NLP4LP dataset, and a less robust LLM agent-based system for modeling optimization problems that we will refer to as OptiMUS-0.2. The new enlarged version of the NLP4LP dataset which accompanies this paper has been open-sourced and includes a modular design that enables future researchers to improve on different sub-problems within our modular optimization framework such as parameter extraction, modeling, or coding. Compared to the earlier version, the modular design of OptiMUS-0.3 substantially improves the robustness of the system for more complicated modeling problems. New modules that enable the system to leverage advanced solver functionality, such as special-order-set constraints, can dramatically speed up the solution time for the associated optimization problems. Additional LLM techniques in OptiMUS-0.3 such as self-reflective error correction improve performance. Finally, this paper is the first to present the OptiMUS webapp.
@@ 2 Related Work
Formulating an optimization model is often a challenging task even for experts in optimization. Different formulations can lead to significantly different solving times and enable the use of different solvers or solution techniques (Boyd and Vandenberghe 2004). One important skill for an optimization expert is to identify assumptions or relaxations that allow for casting the problem as a well-studied problem type, such as MILP, which enables the use of well-developed optimization solvers. Crafting such an efficient formulation often requires specialized knowledge (Zohrizadeh et al. 2020, Low 2013, Roubíček 2020, Luo et al. 2010, Krarup and Pruzan 1983).
@@
Recent progress in Natural Language Processing (NLP) has led to the development of large language models (LLMs) useful for tasks such as answering questions, summarizing text, translating languages, and coding (OpenAI 2023, Touvron et al. 2023, Chowdhery et al. 2022, Wei et al. 2023, Gao et al. 2023a, Borgeaud et al. 2022). Connections to other software tools extend the reach and accuracy of LLMs, as demonstrated by plug-ins for code writing and execution (Paranjape et al. 2023, Wei et al. 2023). Yang et al. (2023) use LLMs to directly generate solutions to optimization problems without calling traditional solvers through prompt optimization to improve performance. The approach is limited to small problems since the performance of LLMs degrades as the input context grows, even for explicitly long-context models (Liu et al. 2023).
@@
In a recent paper, Chen et al. (2023) developed a chatbot to help users detect and fix infeasible optimization problems expressed in Pyomo code and servers as an AI assistant rather than as a solver. Li et al. (2023) designed a chatbot to answer natural-language queries about a supply chain optimization model. Alibaba Cloud (2022) also developed a chatbot to facilitate optimization modeling, but there is no public paper or documentation available on it. Lawless et al. (2024) explored the use of LLMs to allow users to customize a simple constraint programming model in the context of meeting scheduling, but does not model more general MILP optimization problems.
@@
More closely related to our approach, Ramamonjison et al (2023) introduced a dataset of 1101 natural language representations of LP problems. They proposed a two-stage mapping from the natural-language representation to the problem formulation using an intermediate representation. Ramamonjison et al (2022) designed a system to simplify and improve the modeling experience for operations research, but did not offer an end-to-end solution. Xiao et al. (2023) presented a multi-agent cooperative framework called Chain of Experts (CoE) to automatically model and program complex operation research (OR) problems, and evaluated it on NL4Opt and another more complex dataset, ComplexOR, introduced in that paper. Astorga et al. (2024) present a Monte-Carlo Tree Search (MTCS) based approach that decomposes the modeling problem into stages (e.g., modeling variables, then objectives, then constraints, etc.) and uses MTCS to explore the space of plausible models. As mentioned in the paper, this approach is not directly comparable to OptiMUS as it outputs a set of functionally distinct models as opposed to a single formulation and answer. Concurrent with the preparation of this paper, Tang et al. (2024) introduced a new approach to OR modeling with LLMs that uses fine-tuning on a semi-synthetic dataset rather than prompt optimization or agentic models, which they call ORLM (Operations Research Language Model). Yang et al. (2024) also explore a synthetic data generation and fine-tuning approach for a broader class of optimization problems. These fine-tuning approaches are complementary to our work, as any progress towards more capable LLMs for optimization modeling can be combined with our modular prompting framework.
@@ 3 Dataset
As part of this paper, we have curated a comprehensive open-source dataset of 355 optimization problems we call NLP4LP. Our goal in creating NLP4LP is to provide the community with examples they can use to design optimization modeling tools using natural language systems and to assess their quality. The problems in NLP4LP are partitioned into 1) an easy dataset that contains only LP problems with short descriptions and scalar parameters, and 2) a hard dataset that contains both LP and MILP problems with longer descriptions and multi-dimensional parameters (Table 1). Unlike other modeling datasets, each instance in the NLP4LP dataset is accompanied by associated code to run the instance, and ground-truth intermediary representations of the problem (i.e., extracted parameters and targets, a list of clauses of the problem represented in natural language, , and code, and a solution and optimal value for given problem data). This additional structure enables future researchers to investigate and improve different modules of the OptiMUS framework such as parameter extraction or code implementation, and even test these improvements with the OptiMUS webapp, without implementing the full software stack from scratch. Compared to the initial version of the NLP4LP dataset presented in (Anonymous 2024) this version contains more instances from a wider breadth of domains and more complex problems from published papers, and includes additional information for each instance, such as code, as discussed earlier. The dataset is available on Hugging Face and in the supplementary material.
Given the complexity of the task, and the components of each instance (description, sample feasible data, and solution code), gathering high-quality optimization instances is expensive. Unlike traditional ML datasets, NLP4LP does not provide a training set with the same distribution as the test set. As the cost of collecting each instance is high, we have prioritized the task of benchmarking by emphasizing the test set rather than reducing the size of the test set to provide training data. Consequently, NLP4LP is partitioned into 1) a development (dev) set of 22 problems (12 easy and 10 hard) and 2) a test set of 332 problems (277 easy and 55 hard). We designed the dev/test partition intending that the dev set be used to develop optimization modeling tools, while the test set be used to evaluate such tools. In our experiments and system design, we did not use any part of the test set for development, ensuring an unbiased evaluation of the models.
@@ 4 Methodology
A feasible point $x^{*}$ that minimizes the objective function is called a (optimal) solution and has associated optimal objective value $c^{T}x^{*}$. For brevity, we refer to the equations defining constraints and the objective as different clauses in the optimization model. A parameter of an optimization model is any numeric coefficient in the optimization model (i.e., elements of $A$ or $c$). The goal of the OptiMUS system is to go from a natural language description of an optimization problem (e.g., a paragraph of text describing a business problem), which we call the problem description, to a decision (assigment of numerical values to each variable) that solves the problem.
OptiMUS-0.3 decomposes these two tasks into a sequence of LLM-powered steps that incrementally constructs an optimization model and implements it. An overview of the complete workflow of OptiMUS-0.3 is outlined in Fig. 1. At a high level, OptiMUS-0.3 begins by 1) identifying parameters from the problem description and writing a natural-language description of each clause in the problem. It then 2) formulates both the clauses and decision variables of the optimization model with mathematical precision, using  code. Finally, 3) OptiMUS-0.3 uses the  model of the optimization problem and the extracted parameters to generate code snippets for each clause in python and assembles the snippets into a single runnable code file. By running this file, OptiMUS-0.3 outputs an optimal solution and optimal objective value. Throughout this process, OptiMUS-0.3 maintains a state that documents what is known about the optimization model, as well as a connection graph that tracks which variables appear in which constraint. Each LLM-powered component of OptiMUS-0.3 also incorporates an Error Checking (EC) module that catches common errors and improves the accuracy of the modeling and implementation.
@@
The remainder of this section details the key components of the OptiMUS-0.3 framework. Section 4.1 describes the underlying state and connection graph of the system. Section 4.2 describes a sample LLM component of the system. The remaining sections describe specialized modules that improve the performance of the system, including error correction and debugging (Section 4.3), specialized modeling tools (Section 4.4), and advanced optimization coding (Section 4.5). The components described in this Section have been implemented and released as open-source code111included in the supplementary material and are publicly available to use through a web application, described in Section 6.
@@ 4.2 Sample LLM Component
For a comprehensive list of prompts used in OptiMUS-0.3 see our code222included in supplementary materials. We also experimented with incorporating Retrieval Augmented Generation (RAG) into OptiMUS. RAG chooses in-context learning examples that are considered most likely to be helpful for a given challenge, often by choosing examples that minimize some distance metric to the input task. However, RAG had inconsistent impact on performance (see Appendix 10) and is not used in the final version of OptiMUS-0.3.
@@ 4.3.1 Reflective prompts
LLMs can often identify and fix their mistakes by using reflective prompts (Shinn et al. 2023), a strategy by which an LLM is asked to reflect upon its output and decide if it needs to be corrected. We analyzed the most common types of mistakes made by LLMs on the development dataset at each step of the process and designed reflective prompts to address each one (see Appendix 9 for a full list of reflextive prompts used in OptiMUS-0.3). Using these prompts significantly reduces the modeling error rate. For instance, Figure 3 shows an example where the reflexive prompt Are units the same for both sides of this constraint? can help correct a modeling error. The resulting improvement to OptiMUS is substantial (see Table 3), and substantiates the claim that domain-specific knowledge can be used to develop LLM applications that outperform general-purpose LLMs.
@@ 4.3.2 Confidence-based user/LLM feedback
In the context of real-world optimization modeling, it is exceeding rare for a natural-language problem description to correspond to an unambiguous MILP formulation. Rather, the process of optimization modeling generally involves many cycles of formulation and user feedback to arrive at a correct model for the problem. OptiMUS-0.3 simulates this process in order to detect errors and improve its solve rate. OptiMUS-0.3 can use a confidence-based reflective method to decide when to request help with its formulation, either from a user (if available) or another (more powerful) LLM.
First, following the methodology of existing work on confidence calibration in LLMs (Virk et al. 2024), we prompt the system to asses its confidence in its output on a scale of 1 to 5. If the confidence score is less than 4, OptiMUS-0.3 asks the user (or a stronger LLM) to help. The user can decide to either keep the output (e.g., constraint, code), remove it, or modify it. In the OptiMUS webapp (Section 6), this feature is used to highlight outputs that merit particular attention. In our experiments (Section 5), this feature is turned off by default to facilitate streamlined and reproducible testing without a user in the loop. See Fig. 4 for an example of how a weaker LLM (Llama-3-70B) can call on a stronger LLM (GPT-4o) when it is not confident in its solution.
@@ 4.3.3 Debug Code
4.4 Structure Detector Agent
OptiMUS maintains a pool of optimization structures commonly used in optimization software. These structures can be specified explicitly in many modern optimization solver interfaces (Achterberg 2019, CPLEX User’s Manual 1987, Gamrath et al. 2016), which can enhance problem-solving performance and simplify the code. Typical examples of structure include Special Ordered Set (SOS) (Beale and Forrest 1976), indicator variables, semi-continuous variables, and piecewise-linear constraints. At least one of these structures appear in approximately 10% of the NLP4LP dataset (see Appendix 11). We expect these structures to be even more common, and thus important to exploit, in industrial applications.
To detect structure in a constraint (or variable), OptiMUS iterates through these structures and formats them into a structure detection prompt. Within each prompt, the LLM is provided with the structure description, explained by an example illustrating how the structure should be exploited. The LLM is asked to decide whether the structure is relevant to the existing formulation. Upon identifying the appropriate structure, the formulation is adjusted to highlight the problem structure. For instance, a set of constraints that indicate only one of a group of decision variables can be non-zero may be reformulated as type-1 SOS constraint. This structure is then conveyed to gurobi via the python interfaces for special structures such as SOS constraints or indicator variables. This information can either be exploited within the branch and bound solver (e.g., using SOS constraints for branching rules) or allows the solver to automatically reformulate the structure into linear constraints (i.e., a big-M formulation for indicator constraints) (Gurobi Optimization 2023). While it may seem counter-intuitive to identify structure in a MILP only to have that structure re-formulated into linear constraints, in practice we found that automated methods within Gurobi can generate tighter formulations than OptiMUS alone (e.g., a better selection for big-M values). Figure 5 shows two instances where identifying SOS constraints and indicator variables leads to faster performance than a naive implementation. For full details of these examples are included in Appendix 11. The structure detection agent is run during the formulate clauses stage of the algorithm.

### v3

@@
Large Language Models, Optimization Modeling, AI Agents, Mixed Integer Linear Programming, Optimization Tools, Human-Computer Interaction
@@ 1 Introduction
  • We develop a publicly available web application that allows practitioners to try OptiMUS. This webapp enables human-in-the-loop automated optimization modeling: users can edit LLM outputs and OptiMUS can flag low-confidence LLM outputs to solicit corrections from the user.
An initial version of this work was published in a conference proceeding (AhmadiTeshnizi et al. 2024), that introduced a smaller version of the NLP4LP dataset, and a less robust LLM agent-based system for modeling optimization problems that we will refer to as OptiMUS-0.2. The new enlarged version of the NLP4LP dataset which accompanies this paper has been open-sourced and includes a modular design that enables future researchers to improve on different sub-problems within our framework such as parameter extraction, modeling, or coding. Compared to the earlier version, the modular design of OptiMUS-0.3 substantially improves the robustness of the system for more complicated modeling problems. New modules that enable the system to leverage advanced solver functionality, such as special-order-set constraints, can dramatically speed up the solution time for the associated optimization problems. Additional LLM techniques in OptiMUS-0.3 such as self-reflective error correction improve performance. Finally, this paper is the first to present the OptiMUS webapp.
@@ 2 Related Work
Formulating an optimization model is often a challenging task even for experts in optimization. Different formulations can lead to significantly different solving times and enable the use of different solvers or solution techniques (Boyd and Vandenberghe 2004). One important skill for an optimization expert is to identify assumptions or relaxations that allow for casting the problem as a well-studied problem type, such as MILP, which enables the use of well-developed optimization solvers. Crafting such an efficient formulation often requires specialized knowledge (Zohrizadeh et al. 2020, Low 2013, Roub´ıček 2020, Luo et al. 2010, Krarup and Pruzan 1983).
@@
Recent progress in Natural Language Processing (NLP) has led to the development of LLMs useful for tasks such as answering questions, summarizing text, translating languages, and coding (OpenAI 2023, Touvron et al. 2023, Chowdhery et al. 2022, Wei et al. 2023, Gao et al. 2023, Borgeaud et al. 2022). Connections to other software tools extend the reach and accuracy of LLMs, as demonstrated by plug-ins for code writing and execution (Paranjape et al. 2023, Wei et al. 2023). Yang et al. (2023) use LLMs to directly generate solutions to optimization problems without calling traditional solvers through prompt optimization to improve performance. The approach is limited to small problems since the performance of LLMs degrades as the input context grows, even for explicitly long-context models (Liu et al. 2023).
@@
In a recent paper, Chen et al. (2023) developed a chatbot to help users detect and fix infeasible optimization problems expressed in Pyomo code and serves as an AI assistant rather than as a solver. Li et al. (2023) designed a chatbot to answer natural-language queries about a supply chain optimization model. Alibaba Cloud (2022) also developed a chatbot to facilitate optimization modeling, but there is no public paper or documentation available on it. Lawless et al. (2024) explored the use of LLMs to allow users to customize a simple constraint programming model in the context of meeting scheduling, but does not model more general MILP optimization problems.
@@
More closely related to our approach, Ramamonjison et al (2023) introduced a dataset of 1101 natural language representations of LP problems. They proposed a two-stage mapping from the natural-language representation to the problem formulation using an intermediate representation. Ramamonjison et al (2022) designed a system to simplify and improve the modeling experience for operations research, but did not offer an end-to-end solution. Xiao et al. (2023) presented a multi-agent cooperative framework called Chain of Experts (CoE) to automatically model and program complex operation research (OR) problems, and evaluated it on NL4Opt and another more complex dataset, ComplexOR, introduced in that paper. Astorga et al. (2024) present a Monte-Carlo Tree Search (MCTS) based approach that decomposes the modeling problem into stages (e.g., modeling variables, then objectives, then constraints, etc.) and uses MCTS to explore the space of plausible models. As mentioned in the paper, this approach is not directly comparable to OptiMUS as it outputs a set of functionally distinct models as opposed to a single formulation and answer. Concurrent with the preparation of this paper, Tang et al. (2024) introduced a new approach to OR modeling with LLMs that uses fine-tuning on a semi-synthetic dataset rather than prompt optimization or agentic models, which they call ORLM (Operations Research Language Model). Yang et al. (2024) also explore a synthetic data generation and fine-tuning approach for a broader class of optimization problems. These fine-tuning approaches are complementary to our work, as any progress towards more capable LLMs for optimization modeling can be combined with our modular prompting framework.
@@ 3 Dataset
As part of this paper, we introduce a comprehensive open-source dataset of 355 optimization problems we call NLP4LP. Our goal in creating NLP4LP is to provide the community with examples they can use to design optimization modeling tools using natural language systems and to assess their quality. The problems in NLP4LP are partitioned into 1) an easy dataset that contains only LP problems with short descriptions and scalar parameters, and 2) a hard dataset that contains both LP and MILP problems with longer descriptions and multi-dimensional parameters (Table 1). Unlike other modeling datasets, each instance in the NLP4LP dataset is accompanied by associated code to run the instance, and ground-truth intermediary representations of the problem (i.e., extracted parameters and targets, a list of clauses of the problem represented in natural language, , and code, and a solution and optimal value for given problem data). This additional structure enables future researchers to investigate and improve different modules of the OptiMUS framework such as parameter extraction or code implementation, and even test these improvements with the OptiMUS webapp, without implementing the full software stack from scratch. Compared to the initial version of the NLP4LP dataset presented in AhmadiTeshnizi et al. (2024), this version contains more instances from a wider breadth of domains and more complex problems from published papers, and includes additional information for each instance, such as code, as discussed earlier. The dataset is available on Hugging Face and in the supplementary material.
Given the complexity of the task, and the components of each instance (description, sample feasible data, and solution code), gathering high-quality optimization instances is expensive. Unlike traditional ML datasets, NLP4LP does not provide a training set with the same distribution as the test set. As the cost of collecting each instance is high, we prioritize creating a large test set rather than reducing the size of the test set to provide training data. Consequently, NLP4LP is partitioned into 1) a development (dev) set of 22 problems (12 easy and 10 hard) and 2) a test set of 332 problems (277 easy and 55 hard). The dev set is intended to be used to develop optimization modeling tools, while the test set should be used to evaluate such tools. In our experiments and system design, we did not use any part of the test set for development, ensuring an unbiased evaluation of the models.
@@ 4 Methodology
A feasible point $x^{*}$ that minimizes the objective function is called a (optimal) solution and has associated optimal objective value $c^{T}x^{*}$. For brevity, we refer to the equations defining constraints and the objective as different clauses in the optimization model. A parameter of an optimization model is any numeric coefficient in the optimization model (i.e., elements of $A$ or $c$). The goal of the OptiMUS system is to go from a natural language description of an optimization problem (e.g., a paragraph of text describing a business problem), which we call the problem description, to a decision (assignment of numerical values to each variable) that solves the problem. We assume that the problem description contains all the information needed to model the optimization problem. In practice, users may give ambiguous or incomplete descriptions of an optimization problem (see Wasserkrug et al. (2024), Lawless et al. (2024) for a discussion). Helping users to resolve these ambiguities automatically is an exciting direction for future work.
OptiMUS-0.3 decomposes these two tasks into a sequence of LLM-powered steps that incrementally constructs an optimization model and implements it. An overview of the complete workflow of OptiMUS-0.3 is outlined in Fig. 1. At a high level, OptiMUS-0.3 begins by 1) identifying parameters from the problem description and writing a natural-language description of each clause in the problem. It then 2) formulates both the clauses and decision variables of the optimization model with mathematical precision, using  code. Finally, 3) OptiMUS-0.3 uses the  model of the optimization problem and the extracted parameters to generate code snippets for each clause in Python and assembles the snippets into a single runnable code file. By running this file, OptiMUS-0.3 outputs an optimal solution and optimal objective value. Throughout this process, OptiMUS-0.3 maintains a state that documents what is known about the optimization model, as well as a connection graph that tracks which variables appear in which constraint (detailed in Section 4.1). The current state is supplied in the prompt of every LLM-powered component (detailed in Section 4.2). Each LLM-powered component of OptiMUS-0.3 also incorporates an Error Correction (EC) module (detailed in Section 4.3) that catches common errors and improves the accuracy of the modeling and implementation.
@@
OptiMUS-0.3 can be run as an end-to-end pipeline or through an interactive web application designed to facilitate human-system collaboration. The OptiMUS web application uses the same underlying components as the end-to-end framework but provides additional opportunities for users to audit and edit the outputs of each stage of the OptiMUS pipeline. We defer a deeper discussion of the system to Section 6, and mention specific interaction mechanisms available in the user interface earlier when relevant.
The remainder of this section details the key components of the OptiMUS-0.3 framework. Section 4.1 describes the underlying state and connection graph of the system. Section 4.2 describes a sample LLM component of the system. The remaining sections describe specialized modules that improve the performance of the system, including error correction and debugging (Section 4.3), specialized modeling tools (Section 4.4), and advanced optimization coding (Section 4.5). OptiMUS-0.3, and all associated components, have been released as open-source codeand are publicly available to use through a web application, described in Section 6.
@@ 4.2 Sample LLM Component
For a comprehensive list of prompts used in OptiMUS-0.3 see our code111included in supplementary materials. Within the OptiMUS webapp, users can inspect and correct the output of every LLM-powered component of the system.
@@ 4.3.1 Reflective prompts
LLMs can often identify and fix their mistakes by using reflective prompts (Shinn et al. 2023), a strategy by which an LLM is asked to reflect upon its output and decide if it needs to be corrected. We analyzed the most common types of mistakes made by LLMs on the development dataset at each step of the process and designed reflective prompts to address each one (see Appendix 9 for a full list of reflective prompts used in OptiMUS-0.3). Using these prompts significantly reduces the modeling error rate. For instance, Figure 3 shows an example where the reflexive prompt Are units the same for both sides of this constraint? can help correct a modeling error. The resulting improvement to OptiMUS is substantial (see Table 3), and substantiates the claim that domain-specific knowledge can be used to develop LLM applications that outperform general-purpose LLMs.
@@ 4.3.2 Confidence-based user/LLM feedback
In the context of real-world optimization modeling, it is exceedingly rare for a natural-language problem description to correspond to an unambiguous MILP formulation. Rather, the process of optimization modeling generally involves many cycles of formulation and user feedback to arrive at a correct model for the problem. OptiMUS-0.3 simulates this process in order to detect errors and improve its solve rate. OptiMUS-0.3 can use a confidence-based reflective method to decide when to request help with its formulation, either from a user (if available) or another (more powerful) LLM.
First, following the methodology of existing work on confidence calibration in LLMs (Virk et al. 2024), we prompt the system to assess its confidence in its output on a scale of 1 to 5. If the confidence score is less than 5, OptiMUS-0.3 asks the user (or a stronger LLM) to help. The user can decide to either keep the output (e.g., constraint, code), remove it, or modify it. In the OptiMUS webapp (Section 6), this feature is used to highlight outputs that merit particular attention. In our experiments (Section 5), this feature is turned off by default to facilitate streamlined and reproducible testing without a user in the loop. See Fig. 4 for an example of how a weaker LLM (Llama-3-70B) can call on a stronger LLM (GPT-4o) when it is not confident in its solution.
@@ 4.3.3 Debug Code
4.4 Structure Detection Agent
OptiMUS maintains a pool of optimization structures commonly used in optimization software. These structures can be specified explicitly in many modern optimization solver interfaces (Achterberg 2019, CPLEX User’s Manual 1987, Gamrath et al. 2016), which can enhance problem-solving performance and simplify the code. Typical examples of structure include Special Ordered Set (SOS) (Beale and Forrest 1976), indicator variables, semi-continuous variables, and piecewise-linear constraints. At least one of these structures appears in approximately 10% of the NLP4LP dataset (see Appendix 10). We expect these structures to be even more common, and thus important to exploit, in industrial applications.
To detect structure in a constraint (or variable), OptiMUS iterates through these structures and formats them into a structure detection prompt. Within each prompt, the LLM is provided with the structure description, explained by an example illustrating how the structure should be exploited. The LLM is asked to decide whether the structure is relevant to the existing formulation. Upon identifying the appropriate structure, the formulation is adjusted to highlight the problem structure. For instance, a set of constraints that indicate only one of a group of decision variables can be non-zero may be reformulated as type-1 SOS constraint. This structure is then conveyed to gurobi via the Python interfaces for special structures such as SOS constraints or indicator variables. This information can either be exploited within the branch and bound solver (e.g., using SOS constraints for branching rules) or allows the solver to automatically reformulate the structure into linear constraints (i.e., a big-M formulation for indicator constraints) (Gurobi Optimization 2023). While it may seem counter-intuitive to identify structure in a MILP only to have that structure re-formulated into linear constraints, in practice we found that automated methods within Gurobi can generate tighter formulations than OptiMUS alone (e.g., a better selection for big-M values). Figure 5 shows two instances where identifying SOS constraints and indicator variables leads to faster performance than a naive implementation. Full details of these examples are included in Appendix 10. The structure detection agent is run during the formulate clauses stage of the algorithm.